If g(x)= 2x + 5(3x-2)
then what does g(x)=6,x=?

Answer by benazir.sj@gmail.com(109)

g(x)= 2x + 5(3x-2)
given g(x)=6
6=2x+5(3x-2)
6=2x+15x-10
now to find x we have to simplify this
6=2x+15x-10 (Add the 10 over and combine the X's).
6=17x-10
6+10=17x
16=17x
16/17=x
x=16/17